OpenStudy (anonymous):

Solve the equation for y: 4x + 3y = -12

OpenStudy (anonymous):

is any point on the plane given ??

OpenStudy (anonymous):

because there can be many solutions

OpenStudy (kropot72):

Can you subtract 4x From both sides?

OpenStudy (anonymous):

for many values of x u get different y's

OpenStudy (anonymous):

dude kropot it is still the same

OpenStudy (anonymous):

A.y = -12 - 4/3 x B.y = -4 - 4/3 x C.y = -12 + 4x D.y = -4 + 4x Are the choices given

OpenStudy (kropot72):

Can you subtract 4x from both sides?

OpenStudy (anonymous):

The question here is asking you to make \(y\) a function of \(x.\) Do you know how to do that?

OpenStudy (anonymous):

ofcourse i do

OpenStudy (anonymous):

I was talking to @simpleeLizzy.

OpenStudy (anonymous):

I dont

OpenStudy (anonymous):

my bad

OpenStudy (anonymous):

check my ans @simpleeLizzy

OpenStudy (anonymous):

@simpleeLizzy, do you know how to solve equations for a particular variable?

OpenStudy (anonymous):

to be honest no ; math isnt my thing /;

OpenStudy (anonymous):

4x+3y=-12 3y=-12-4x y=-4-4/3x

OpenStudy (anonymous):

Okay, well.. Here's what you can do. You can get the \(x\) on the left side. To do this, you subtract it from both sides. Do you follow so far?

OpenStudy (anonymous):

yhis is it

OpenStudy (anonymous):

this is it

OpenStudy (kropot72):

(4x - 4x) + 3y = -12 - 4x 0 + 3y = -12 - 4x @simpleeLizzt Do you follow?

OpenStudy (anonymous):

check my steps

OpenStudy (anonymous):

@karthik_a.k.a_joker, the point is to show the asker how to get to that answer, not to just outright state it.

OpenStudy (anonymous):

my steps were showing the procedure of how to approach the problem

OpenStudy (anonymous):

Ahhh im just getting more confused

OpenStudy (anonymous):

check my steps @simpleeLizzy

OpenStudy (anonymous):

dnt get confused it is easyyyyyyyy

OpenStudy (anonymous):

@karthik_a.k.a_joker its one of the choices

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thanks to all (;

OpenStudy (anonymous):

y=(-12-4x)/3

OpenStudy (anonymous):

4x + 3y = -12 Subtract 4x from both sides: 3y= -4x-12 Divide both sides by 3: Y=1/3 (-12-4x) Result: Y=-4/3 (x+3)

OpenStudy (anonymous):

I hope this help.

OpenStudy (anonymous):

are u insane it one equation & 2 unknown